# A Review of Community-Centric Power Systems Resilience Assessment and Enhancement Strategies

### v1

@@ Abstract
This paper presents a comprehensive review of resilience metrics, covering both engineering-based measures, such as fragility-curve modeling, and data-driven approaches, including triangular and trapezoidal representations. Next, the paper examines the interdependencies between power systems resilience and community resilience, addressing socioeconomic and behavioral dimensions, infrastructure interconnections, and the emerging role of resilience hubs. The review then synthesizes state-of-the-art strategies for enhancing power system resilience, including network hardening, resource allocation, optimal scheduling, and reconfiguration techniques. Special emphasis is placed on the integration of Artificial Intelligence (AI) methods and the techno-legal dimensions of resilient power systems and communities. In particular, the paper contrasts the regulatory landscapes of the European Union and the United States, highlighting key similarities and distinctions. By analyzing methodologies for mitigating the impacts of high-impact, low-probability (HILP) events, the review identifies critical research gaps and outlines promising directions for future investigation.
@@
Community resilience, high-impact low-probability events, power systems resilience, resilience evaluation methods, resilience enhancement strategies, resilience metrics.
@@ 1 Introduction
The frequency of power outages and their associated losses have been growing in the U.S. over the past 43 years, as illustrated in Figure 1[1]. Aging infrastructure and the rise of extreme events have highlighted the vulnerability of power systems [2]. Storm-related disruptions are estimated to cost the U.S. economy between \$20 billion to \$55 billion annually [3]. Several research studies anticipate the frequency and severity of such extreme events to increase due to climate change [3, 4, 5]. This underscores the critical need to adapt power grids to withstand disasters and mitigate catastrophic impacts. Ensuring the resilience of the energy supply and infrastructure is essential for facilitating economic recovery and fostering sustainable development.
Various definitions, metrics, assessment methods, and improvement strategies for power system resilience have been proposed. However, the absence of standardized definitions and metrics poses a challenge in accurately estimating the resilience of power systems [6]. The U.S. Department of Energy (DOE) defines energy system resilience as “the ability of the grid, buildings, and communities to withstand and rapidly recover from power outages while continuing to provide electricity”. According to the IEEE, resilience is defined as “the ability to withstand and reduce the magnitude and/or duration of disruptive events, which includes the capability to anticipate, adapt to, and/or rapidly recover from such an event.” Community resilience builds upon this concept by extending it to social and infrastructural systems, emphasizing collective preparedness and adaptive capacity which is defined as a community’s ability to reduce and prepare for risks [83]. Berkes et al. [84] proposed two perspectives on community resilience: the social-ecological systems perspective and the psychology and community development perspective. Various socioeconomic and demographic factors that shape community resilience also influence power system resilience [70, 72]. Additionally, power system resilience assessments must consider the interdependency of power systems with other critical infrastructure, as cascading failures can impact multiple systems. Proper parameter tuning is essential for maintaining the stability of integrated networks [81, 82].
Several research studies have extensively analyzed the concepts of resilience and strategies for enhancement, providing a detailed overview encompassing definitions, evaluation methodologies, modeling approaches, and strategies for enhancing power system resilience [9, 10, 11, 12]. Despite these contributions, there is a lack of comprehensive exploration of the resilience triangle and trapezoid frameworks, detailed assessment methodologies, and the integration of modeling techniques with AI and ML for resilience enhancement. Additionally, comprehensive discussions linking policy, regulatory frameworks, and socio-economic factors to resilience enhancement remain limited. This paper addresses these gaps by providing an in-depth analysis of power system resilience evaluation and enhancement methods against extreme weather events, with a primary focus on the state-of-the-art AI and ML techniques.
The structure of the paper is organized as follows. Section 2 explores resilience assessment approaches. Section 3 presents a discussion on the interdependence of power system resilience and community resilience. Section 4 delves into resilience enhancement strategies, then Section 5 presents the application of AI in resilience analysis. Section 6 reviews the techno-legal aspects of resilient power systems and communities. Section 7 identifies research gaps, outlines challenges, and proposes guidance for future research. Finally, Section 8 concludes the paper with a summary of the key findings and a discussion of the identified limitations within current research efforts.
@@
2 Power System Resilience Assessment Practices
Effective implementation of resilient planning and operations depends on a comprehensive evaluation of resilience using standardized metrics. Methods for assessing resilience in power systems have received increasing attention in recent research and can be broadly classified into two main approaches: 1) Detailed engineering (micro-level), and 2) Data-driven (macro-level).
2.1 Detailed Engineering Resilience Assessment
Detailed engineering resilience assessment techniques emphasize analyzing the specific fragility curves of individual components within a system, offering a granular perspective on their vulnerability to extreme events. A fragility curve represents the probability of failure of a system or one of its components under a specific event. In scenarios with minimal uncertainty, the fragility curve closely resembles a step function, transitioning sharply between two states: 0 (non-failure) and 1 (failure) [32]. However, in most real-world applications, fragility curves are modeled using lognormal cumulative distribution functions, which rely on the logarithmic mean and logarithmic standard deviation to provide a more realistic representation by accounting for variations in system behavior under different load conditions. In [34], the fragility curve is described as a lognormal distribution
$f(x)=\frac{l}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{x-\mu}{\sigma})^{2}},$ (1)
where $x$ represents event intensity, $\mu$ and $\sigma$ denote the logarithmic mean and standard deviation, respectively. The slope of the fragility curve is influenced by these two parameters, determining the system’s sensitivity to variations in event intensity. For example, Figure 2 illustrates fragility curves for transmission lines and towers under varying wind speeds (a) and infrastructure failure as a function of peak ground acceleration (PGA) during earthquakes (b). The probability of failure ($P_{f}$) in seismic scenarios can be expressed as:
$P_{f}(PGA)=\alpha\left[\frac{1}{\beta_{tot}}\ln\left(\frac{PGA}{PGA_{mi}}\right)\right],$ (2)
where $P_{f}$ denotes the probability of infrastructure failure, $\alpha$ is the standard cumulative probability function, $PGA_{mi}$ is the median threshold value of the peak ground acceleration, and $\beta_{tot}$ is the total standard deviation [35].
Figure 2: Fragility curve vs extreme event intensity as a function of wind speed (a), and of PGA (b).
Ranjbar et al. [19] analyzed fragility curves for substations, wind farms, and transmission lines during hurricanes. Substation damage probabilities were modeled using log-normal distributions, transmission structure failures with exponential functions, and wind turbine failures with log-logistic distributions based on wind speeds. Four methodologies have been established for constructing fragility curves for power system components: 1) judgmental, based on expert experience, 2) empirical, based on field and laboratory measurements to model the performance of the components [33], 3) analytical, derived from structural simulation models, and 4) hybrid, a combination of the judgmental, empirical, and analytical methods.
Sabouhi et al. [17] proposed a framework for resilience assessment of power systems that provides a step-by-step process for assessing the Resilience Index (RI) using fragility curves. Panteli et al. [33] presented a similar framework to evaluate the resilience of a system using fragility curves and a sequential Monte Carlo Simulation method. Work in [31] introduced a cost-effective approach for system operators to calculate RI using fragility curves. One drawback of the detailed engineering approach is that it requires comprehensive information about the power network and its components, which is generally not publicly available data.
2.2 Data-driven Resilience Assessment
The data-driven resilience assessment approach, commonly known as the system-level approach, employs comprehensive data analytics to assess resilience metrics. It focuses on assessing the performance of the system that influences its ability to withstand and recover from disruptions. Operational resilience describes the capacity of a power system to sustain generation availability and continuous electricity delivery to consumers during and after disruptive events such as disasters, whereas infrastructure resilience focuses on the power system’s physical robustness and minimizing component failures or disruptions. Together, these aspects are reflected in probabilistic resilience curves to characterize system performance across a range of disturbance types, impacts, and severity [24],[8].
2.2.1 Resilience Triangle
The resilience triangle theory was developed by Bruneau et al. [7, 20] and serves as the foundation for resilience assessment. It primarily examines how a system’s performance deteriorates throughout an event. According to Tierney and Bruneau [7, 20], a system exhibits resilience when it minimizes the probabilities of failure and the impacts of failures, including damage, loss of life, and negative social and economic consequences, while also achieving shorter recovery times.
The resilience triangle method, depicted in Figure 3, provides a quantitative assessment of resilience via a time-dependent metric. The resilience metric evaluates system performance on a scale from 0% (a complete loss of service) to 100% (no disruption in service). The targeted performance is represented by $P_{T}$, while actual performance for operational and infrastructure resilience is denoted by $P_{op}(t)$ and $P_{in}(t)$ accordingly. As depicted in Figure 3, an extreme event occurring at time $t_{1}$ may cause substantial damage to the network, leading to a significant reduction in quality when it reaches $P_{op}^{min}$ (for operation) or $P_{in}^{min}$ (for infrastructure) at time $t_{2}$. After the event, recovery progresses until full restoration at $t_{4}$.
Figure 3: Schematic diagram of the resilience triangle.
A time-dependent resilience metric, $RI_{s1}$, is expressed in (3), which evaluates system resilience using both the resilience triangle and trapezoid. This metric is derived from the ratio of the area under the actual performance curve, $P_{op}(t)$, to the area between the target performance curve, $P_{T}$ [21, 87]. This approach can be applied to both operational and infrastructure resilience indexes,
$RI_{s1}=\frac{\int_{t_{1}}^{t_{4}}P_{op}(t)dt}{\int_{t_{1}}^{t_{4}}P_{T}dt}.$ (3)
In addition to the normalized metric presented in (3), Gholami et al. [22] introduced another metric,
$RI_{s2}$ $=\frac{\int_{t_{0}}^{t_{4}}[P_{T}-P_{op}(t)]dt}{\int_{t_{0}}^{T}P_{T}dt}=1-\frac{\int_{t_{0}}^{t_{4}}P_{op}(t)dt}{\int_{t_{0}}^{t_{4}}P_{T}dt}=1-RI_{s1}.$ (4)
In contrast to $RI_{s1}$, $RI_{s2}$ quantifies the reduction in resilience from the ideal value of 1 due to an extreme event, where zero signifies a fully resilient system and one represents a non-resilient system.
Moreover, G. Cimellaro et al. [23] articulated the resilience triangle metric for infrastructure to extreme events as
$RI_{s3}=\int_{t_{1}}^{T}\frac{P_{in}(t)}{T-t_{1}}dt,$ (5)
where
$P_{in}(t)=$ $[1-L(I,T_{RE})][H(t-t_{1})-H(t-(t_{1}+T_{RE}))]$
$\times f_{REC}(t,t_{1},T_{RE}),$ (6)
and
$T_{RE}=$ $t_{1}-t_{4}.$ (7)
Here, $L(I,T_{RE})$ represents the loss function, wherein $I$ and $T_{RE}$ denote the earthquake intensity and recovery time from the event, while $H(.)$ denotes the Heaviside step function. The parameter $T$ stands for the control time of the system after the restoration state. Additionally, $f_{REC}(t,t_{1},T_{RE})$ characterizes the recovery function. The authors emphasized metrics to measure resilience dimensions such as rapidity, redundancy, robustness, and resourcefulness.
While the resilience triangle is highly valued for its applicability across various systems, it lacks the capability to measure other pivotal resilience dimensions encountered by standard power systems. For instance, it fails to capture the duration infrastructure spends in the post-degradation state (from $t_{2}$ to $t_{3}$ in Figure 4) before restoration efforts begin and conclude [8]. To address these limitations, the resilience triangle concept has evolved into the resilience trapezoid framework [10], [13].
2.2.2 Resilience Trapezoid
The resilience trapezoid is the enhanced model of the resilience triangle, which incorporates the system’s degraded state in the event of delayed restoration actions, thereby providing a more realistic and comprehensive approach to understanding and measuring resilience [8],[25]. Illustrated in Figure 4, a conceptual resilience trapezoid delineates the states of a power system in response to an external disruption, alongside the temporal progression of these states over time [15], [25].
Figure 4: Schematic diagram of the resilience trapezoid.
Arjomandi-Nezhad et al. [26], proposed a recovery-based resilience metric which is defined as the ratio of the number of recovered customers to the average outage time of the affected customers
$RI_{s4}=\frac{P_{T}-P_{op}^{min}}{\frac{\int_{P_{op}^{min}}^{P_{T}}tdP}{P_{T}-P_{op}^{min}}},$ (8)
which can be rewritten as
$RI_{s4}=\frac{(P_{T}-P_{op}^{min})^{2}}{(T-t_{2})\times P_{T}-\int_{t_{3}}^{T}P(t)dt-\int_{t_{2}}^{t_{3}}P(t)dt}.$ (9)
In (9), $T$ represents the time at the end of the post-restoration state and $P_{op}^{min}$ represents the minimum level of the performance parameter during an extreme event. Similar to (8) and (9), equivalent formulations can be developed for assessing the resilience of system infrastructure. The model aims to restore as many loads as early as possible by maximizing the area under the recovery curve to achieve a higher resilience value.
Wang et al. [27] assessed $RI$ to enhance power resilience by finding the optimal size of Micro-grids (MGs), considering both cost and resilience factors. The authors introduced a version of $RI$ that denotes the proportion of weighted served loads after an extreme event, as expressed in (10). A smaller $RI$ indicates poor MG performance, and an $RI$ of 1 indicates no load shedding.
$RI_{s5}=\frac{\int_{t_{1}}^{t_{4}}P_{op}(t)dt}{P_{T}\times(t_{4}-t_{1})},$ (10)
where
$P_{T}(t)=w_{c}P_{c}+w_{n}P_{n},$ (11)
$P_{op}(t)=w_{c}P_{c}(t)+w_{n}P_{n}(t).$ (12)
Here, $P_{c}(t)$ and $P_{n}(t)$ indicate the power served to critical and non-critical loads. By assigning higher weights to critical loads, the metric shows that the system can remain resilient even if several non-critical loads are disconnected during the disturbance.

### v2

# A Review of Community-Centric Power System Resilience: Strategies, Data-Driven Methods, and Techno-Legal Perspectives

@@ Abstract
This paper presents a comprehensive review of community-centric power system resilience, emphasizing the integration of community-level resilience considerations and techno-legal governance frameworks with engineering-based resilience enhancement strategies and data-driven approaches to address extreme events. Recent large-scale outages have demonstrated that power disruptions can cascade beyond electrical infrastructure and disproportionately affect vulnerable communities, critical services, and interconnected urban systems, highlighting the need for resilience approaches that integrate technical, social, and regulatory dimensions. Within this community-centric perspective, the review first summarize state-of-the-art strategies for enhancing power system resilience, including network hardening, resource allocation, optimal scheduling, and system reconfiguration techniques, while highlighting the growing role of artificial intelligence (AI) and data-driven analytics in supporting resilience planning and operational decision-making. It then examines the interdependencies between power system resilience and community resilience, addressing socioeconomic and behavioral dimensions, cross-infrastructure interconnections, and the emerging role of resilience hubs. The paper further examines the techno-legal frameworks governing resilient energy systems by comparing the regulatory landscapes of the European Union (EU) and the United States, highlighting key similarities and distinctions that shape resilience planning and implementation. By analyzing state-of-the-art engineering-based, AI-driven, and techno-legal methods for assessing and mitigating the impacts of high-impact, low-probability (HILP) events, the review identifies critical research gaps and outlines promising directions for future investigation.
@@
Community resilience, high-impact low-probability events, power system resilience, resilience enhancement strategies, artificial intelligence in power systems, techno-legal frameworks.
@@ 1 Introduction
The frequency of power outages and their associated losses have been growing in the U.S. over the past 43 years, as illustrated in Fig. 1 [45]. Aging infrastructure, increasing system complexity, and the rising intensity of extreme events have exposed critical vulnerabilities in modern power systems [73]. Storm-related disruptions are estimated to cost the U.S. economy between \$20 billion and \$55 billion annually [23]. Historical events further highlight the societal consequences of prolonged outages. For example, Hurricane Maria in 2017 caused a widespread collapse of the electric power system in Puerto Rico, leaving many communities without electricity for months and contributing to an estimated 2,975 excess deaths [64]. Similarly, the 2021 Texas winter storm (Storm Uri) triggered widespread power outages affecting more than 4.5 million customers disproportionately impacting low-income households and critical community services [97]. More recent large-scale outages, including the April 2025 Iberian Peninsula blackout and the 2025 California blackout, demonstrate how extreme-event-driven power failures can cascade beyond the electric grid, disrupting transportation and communications systems  [101, 33]. Several studies further anticipate that both the frequency and severity of such high-impact events will continue to increase under climate change [23, 82, 92]. This underscores the critical need to adapt power grids to withstand extreme disturbances and mitigate catastrophic impacts.
In this context, power system resilience has emerged as a key framework for guiding the analysis and design of modern power grids. It refers to the ability of the system to withstand, adapt to, and rapidly recover from disruptive events, while also encompassing broader socio-economic dimensions through the concept of community resilience [89].
@@
Power system resilience assessment can generally be categorized into two complementary paradigms: component-level engineering approaches and system-level data-driven approaches. Engineering-based methods rely on fragility modeling of individual components—such as transmission lines, substations, and generation assets—to quantify failure probabilities under extreme events and evaluate physical vulnerability scenarios [81, 86]. While these approaches provide valuable insight into failure mechanisms, their practical application is often constrained by limited access to detailed component-level data. In contrast, data-driven approaches evaluate resilience using aggregate performance metrics—such as resilience curves—to characterize system-level degradation and recovery [20, 80, 58, 19].
Prior studies have shown that community resilience is shaped by multiple socioeconomic and demographic factors, which also influence power system resilience and recovery outcomes [68, 32]. Moreover, power systems are tightly interdependent with other critical infrastructures, including transportation, communication, and healthcare systems, such that failures in one domain can propagate and amplify impacts across others. These interdependencies highlight the importance of coordinated resilience planning to maintain stability in integrated energy systems [128, 12]. Similar ideas are increasingly discussed in the smart infrastructure and smart city literature, where resilience is framed as a socio-technical challenge [3].
A substantial body of literature has reviewed resilience concepts, assessment frameworks, and enhancement strategies for power systems [4, 58, 16, 71]. However, most existing reviews primarily focus on technical resilience enhancement methods or analytical assessment approaches, while comparatively limited attention has been given to integrating these strategies with community vulnerability considerations and the techno-legal frameworks that govern resilience planning and implementation. As power grids become increasingly digitized and interconnected with other critical infrastructures, resilience planning must address not only engineering reliability but also social equity, cyber-physical risks, and regulatory compliance. This gap highlights the need for a comprehensive review that bridges these technical, social, and regulatory dimensions within a unified community-centric framework.
Motivated by these challenges, this paper presents a comprehensive review of community-centric resilience enhancement strategies, with a focus on critical infrastructure interdependencies, AI- and ML-enabled approaches, and associated regulatory frameworks. This integrated perspective provides a holistic synthesis for understanding how technical resilience strategies interact with societal impacts and regulatory environments in modern energy systems.
The remainder of the paper is organized as follows. Section 2 reviews resilience enhancement strategies for power systems under extreme events. Section 3 presents recent advances in applying AI and ML techniques to resilience analysis and enhancement. Section 4 discusses the interdependence between power system resilience and community resilience. Section 5 reviews techno-legal and regulatory considerations relevant to resilient power systems and communities. Finally, Section 6 summarizes the main findings, identifies research gaps, discusses limitations and opportunities for advancing community-centric power system resilience, and proposes directions for future work.
2 Power Systems Resilience Enhancement Strategies
Resilience assessment provides quantitative methods to identify vulnerabilities of power systems, enabling proactive strategies to strengthen resilience before extreme events occur. Here, different types of resilience-enhancing strategies are identified and explained.
2.1 Hardening Power Network
System hardening refers to modifications made to power system infrastructure to minimize vulnerability to severe events [71, 18]. Hardening is particularly effective at mitigating the impact of extreme events in distribution systems. Typical hardening techniques include reinforcing substations, upgrading or undergrounding power lines, and armoring poles [114]. The U.S. National Electric Safety Code requires the replacement of poles when their structural ability falls below two-thirds of their original strength [114].
In [67], a tri-level optimization is proposed that focuses on hardening power networks to enhance the resilience of distribution systems. It explores strategic decision-making for system hardening through scenario planning under maximum-damage conditions. System hardening, while serving as an effective resilience enhancement strategy, involves a significant financial burden. Consequently, substantial research has focused on developing cost-effective hardening solutions without compromising safety or efficiency [71]. In addition to general infrastructure reinforcement, several studies have examined hardening strategies specifically aimed at mitigating lightning-induced failures in transmission systems operating in regions with high lightning density. Lightning strikes are a major cause of transmission line outages. To reduce these risks, mitigation measures such as improving tower-footing grounding resistance, installing additional shield wires or underbuilt ground wires, and deploying transmission-line surge arresters (TLSAs) have been widely investigated. Recent studies have proposed optimization-based frameworks to determine the optimal placement of TLSAs along transmission corridors. For example, several works validated these strategies using realistic transmission-line models in lightning-prone regions such as Brazil and Malaysia, demonstrating significant reductions in lightning-induced outage rates when protection devices are strategically deployed [112, 111, 110, 25, 5]. In parallel, risk-based resilience frameworks have been developed to quantify the probabilistic impact of thunderstorms on transmission infrastructure by combining lightning detection data, line outage records, and fragility models. These approaches enable utilities to prioritize protection upgrades and maintenance actions based on lightning exposure and asset vulnerability, with applications demonstrated using transmission outage data from Southwest China and probabilistic analyses calibrated to England’s transmission networks [15, 99]. Beyond these infrastructure-level mitigation strategies, the integration of distributed energy resources (DERs) into modern power systems introduces additional challenges. The uncertainty associated with renewable energy operation complicates traditional scenario-based hardening techniques, as it increases the number and severity of system contingencies and potential worst-case scenarios [114].
2.2 Resource Allocation
Resource allocation is a resilience enhancement approach that involves strategically deploying DERs at key nodes in the power system to bolster resilience [9]. Given the uncertainty of DERs operation during extreme weather conditions, it is essential to back up the lost power to maintain system stability. This helps strengthen grid resilience while maintaining reliable operation during both normal and extreme conditions [108]. Strategically positioning DERs within the network can reduce the need for critical load curtailment and enhance overall system resilience [30]. In [123], a two-stage stochastic mixed-integer programming model was developed to optimize the location and sizing of DERs (in the first stage), minimize the operating costs of DERs and load shedding (in the second stage). Other design decisions, like optimal DER size, can be made using similar optimizations with the objective to maintain resilience above a specified threshold [118]. While optimization-based approaches can determine the optimal placement and sizing of DERs under predefined conditions, the machine learning methods reviewed in Section 3 provide data-driven support for resource allocation by leveraging the increasing availability of large-scale data from monitoring systems and intelligent devices [120].
2.3 Optimal Scheduling of Repair Crews
After an extreme incident, quick repairs of the power system are needed to minimize load-shedding losses. Maintaining resilience requires a proactive repair crew deployment approach to prepare for post-disaster repair and restoration. Optimization models have been proposed to minimize load shedding by sending repair crews to the ideal locations before extreme events [17]. Their goal is to maximize the expected restored load while minimizing the repair duration. In [10], Anmar Arif et al. propose a MILP based co-optimization of repairs, reconfiguration, and DER dispatch in a two-stage method for outage management. However, these optimization problems operate under assumptions of full system observability, known damage states, and possible actions. As reviewed in Section 3, deep reinforcement learning methods in references [59] and [130] learn adaptive policies from state–action–reward interactions that generalize across system states and scale to large networks without relying on predefined scenario sets.
2.4 Network Reconfiguration
Network reconfiguration refers to the modification of the network topology by opening normally closed sectionalizing switches and closing normally open tie switches. In order to follow a set of operation constraints, network reconfiguration problems are typically modeled using various objectives, like minimum switching operations, operation costs, or power loss [98]. A number of technologies are employed to improve reconfiguration capacity and distribution system resilience, including smart transformers, DERs, fault prevention and detection systems [71]. Network reconfiguration can be particularly effective at improving power system resilience when more sophisticated measuring instruments are used. Devices like distributed remote controllers, phasor measurement units (PMUs), communicating relays, reclosers, advanced supervisory control and data acquisition (SCADA), and outage management systems (OMS) allow for remote reconfiguration, which can reduce load interruption times. A more effective approach uses the real-time status of the system during each stage of an event to guide subsequent decisions [114]. In Section 3, it will be discussed that AI and machine learning methods enable more adaptive and real-time network reconfiguration by learning control policies that map system states to switching actions.
2.5 Protection Strategies for Resilient Grid Operation
Protection systems represent a critical operational layer for maintaining resilient grid performance during faults and extreme disturbances. Traditional protection schemes rely on fixed relay settings designed for relatively stable operating conditions. However, increasing penetration of DERs, bidirectional power flows, and dynamic network configurations have reduced the effectiveness of conventional protection coordination [1]. Consequently, recent research has focused on adaptive and communication-assisted protection strategies capable of responding to real-time system conditions and limiting disturbance propagation.
One emerging direction involves adaptive relay coordination, where protection settings are dynamically adjusted to accommodate changing operating conditions. For instance, a dual-setting protection strategy for microgrids integrates directional overcurrent relays with communication-assisted coordination to improve relay selectivity and response time [6]. The proposed scheme demonstrated substantial improvements in fault isolation performance. Similarly, adaptive relay frameworks have been proposed to modify the time-multiplier settings of overcurrent relays in real time based on fault impedance variations and operating conditions [113].
Beyond local relay adaptation, wide-area protection systems have been investigated as a complementary mechanism for improving system-level resilience. These approaches utilize synchronized measurements from PMUs to detect faults and evaluate protection system performance across multiple network locations [125]. Simulation studies showed that the method can detect faults within approximately 100 ms and rapidly determine whether primary protection has operated correctly.
Another important class of resilience-oriented protection strategies involves adaptive remedial action schemes (RAS). Unlike conventional RAS implementations with fixed activation conditions, recent studies have explored optimization-based frameworks capable of updating RAS triggers and corrective actions in real time. Optimization-driven RAS design has been shown to significantly reduce post-contingency constraint violations by considering multiple operating scenarios and dynamically adjusting corrective actions under changing system conditions [85].
Communication-enabled protection strategies have also gained attention as a cost-effective method for improving protection selectivity in DER-dominated distribution networks. A recent study introduced a hybrid overcurrent protection scheme that integrates conventional relays with communication-based angle comparison logic [14].
The resilience enhancement strategies presented in this section can be understood within a unified temporal framework of system response to extreme events. Fig. 2 shows the extreme event lifecycle and positions each strategy according to its operational role across pre-event preparedness, event onset, degradation, and recovery phases. Hardening measures are exclusively pre-event investments, while resource allocation spans the pre-event and onset phases; protection strategies activate at fault onset to limit degradation; and network reconfiguration and repair crew scheduling operate continuously from event onset through full restoration.
Figure 2: Power system resilience enhancement strategies to the extreme event lifecycle.
3 AI Methods for Power Systems Resilience
AI and ML methods are emerging as powerful power system operation and planning tools applied to a variety of tasks, particularly to improve the effectiveness of the resilience enhancement methods discussed in Section 2 [42, 43, 120, 107, 63]. Fig. 3 illustrates the integration of heterogeneous datasets into analytics modules for spatiotemporal modeling, prediction, and optimization, enabling decision support such as resilience enhancement.
Figure 3: AI-enabled decision support framework for power system resilience.
Most ML techniques can be categorized into four groups: 1) Traditional ML, 2) Artificial Neural Network (ANN), 3) Reinforcement Learning (RL), and 4) Advanced Generative and Language Models. Table 1 summarizes different ML approaches for power system resilience, highlighting their performance, limitations, and related extreme events. The variation in reported accuracy across studies is primarily attributed to differences in evaluation metrics, nature of considering extreme events, and prediction horizon. High accuracy (e.g., >>90%) often results from simplified classification tasks (0-no outage, 1-outage), or event-specific (e.g., hurricane, lightning) outage prediction while lower accuracy (<<80%) typically reflect longer prediction horizons, and high resolution (e.g., hourly, daily).
Power outage prediction has emerged as a critical area of research to bolster the resilience of energy systems against various disruptions, including weather events, equipment malfunctions, and grid instabilities. Logistic regression is a simple, fast, and robust method for outage prediction [34]. Higher-order and weighted logistic regression improve performance by capturing nonlinear relationships. Some techniques offer higher accuracy than logistic regression. For instance, Support Vector Machines (SVM) has been utilized to predict outages by assessing the condition of grid components [35]. One other notable method is to use ensemble learning techniques. For example, ADABOOST+ has been used to analyze the impacts of lightning and wind on power outages across four cities in Kansas [61]. In [122], the authors enhance prediction model accuracy by pre-classifying events into three severity levels using Quantile Weight Distance (QWD): low (<< 100 outages), moderate (100–1,000 outages), and high (>> 1,000 outages). A conditioned outage prediction model (OPM) is then developed using a Bayesian Additive Regression Tree (BART), Random Forest (RF), and Ensemble Decision Tree (ENS) to outperforms the standard OPM.
While traditional machine learning models are limited in capturing complex nonlinear relationships, ANNs address this limitation, with additional hidden layers. In [104], Convolutional Neural Networks (CNN) combined with long-short-term memory (LSTM) have been applied to predict outages across selected counties in New York, studying the socioeconomic impacts on vulnerable populations based on household income and racial demographics. In many cases, model performance is improved through effective data preprocessing, such as synthetic minority over-sampling technique (SMOTE) [119]. In [116], the role of socioeconomic variables in power outage prediction is also examined.
Category Methodology Key Findings Drawbacks Extreme Event
Advanced Generative and Language Models Transformer-based [57, 124] A type of neural network architecture that process data in parallel rather than sequentially. Reduce forecast errors compared with numerical models. Dependence on multi-source data availability and integration Natural disaster, cyber-attack.
LLM-based Models [129, 7, 124] An application built on Transformer architecture, accuracy can be improved through parameter tuning and contextual adaptability. High computational and storage cost Hurricanes, cyber-attacks, natural disasters.
Reinforcement learning (RL) has emerged as a prominent technique in stability control and restoration tasks. In RL applications, the environment is modeled as a finite-horizon Markov decision process, defined by the tuple $(S,A,p,r,d_{0},\gamma)$. Here, $S$ and $A$ represent the state and action spaces, respectively. $p(s^{\prime}\mid s,a)$ is the state transition function indicating the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$; $r(s,a)$ is the reward function for taking action $a$ in state $s$; $\gamma$ is the discount factor for future rewards, and $d_{0}$ is the initial state distribution [28].
Different RL tasks are characterized by unique state and action spaces, along with task-specific reward functions focused on using RL for microgrid (MG) formation post-disaster [130, 49, 2, 48, 102]. These models often have discrete action spaces (e.g., toggling switches on and off for specific DERs), and sometimes continuous spaces (e.g., adjusting setpoints for selected DERs) [2, 28]. In addition, RL has been applied to optimize restoration strategies, such as dispatching mobile energy resources, and managing repair crews [50, 117]. Another study discuss scheduling repairs for components and compares the performance of the Q-learning algorithm with Greedy Search (GS) and Exhaustive Search (ES) [66].
Other ML approaches are also instrumental in enhancing power system resilience. Predictive control policies for MG management have employed tree-based models to forecast load demands and solar production [55]. Bayesian Networks and hybrid statistical algorithms like the Fragility-curve Monte Carlo Simulation Scenario Reduction (FC-MC-SCENRED) have been compared for pre-hurricane system optimization for distribution networks [77]. Graph neural networks (GNNs) have been proposed to assess node criticality and identify vulnerable lines in cascading outage studies [52, 83]. For disturbance management, CNNs paired with autoencoders have been developed to dynamically adjust shunt capacitors and transformers during electromagnetic pulse disturbances [126]. A one-dimensional CNN has been designed to detect data intrusions in smart meters [121]. Advanced deep learning techniques have been applied for system-wide online load restoration with high wind power penetration [130], and MG planning in active distribution systems using Bayesian Regularization Backpropagation [109].
Recent advancements in transformer-based generative intelligence offer potential to apply in the power systems to better capture complex patterns between parameters. The transformer-based model can tune millions of parameters at a time [124]. Such large-scale data-driven modeling capabilities can also support emerging digital twin frameworks for community-scale energy systems, where integrated models simulate infrastructure behavior, outage impacts, and resilience interventions across interconnected urban systems [96]. The models, such as eGridGPT, are used to train and forecast more accurate power grid analysis and decision making [27]. Domain-specific large language model (LLM) framework extends beyond conventional predictive learning by introducing reasoning, context awareness, and multi-modal understanding of the grid’s characteristics [124]. Self-attention mechanisms can effectively combine information from different sources and understand how events change over time and locations [57, 7]. Another model predicts outage occurrence as a binary classification problem and is evaluated using Precision, Recall and Area Under the Precision-Recall Curve (AU-PRC), enabling to predict up to several hour ahead of the event [7]. However, the model requires high-performance computing infrastructure to process datasets for training and real-time inference.
Despite advantages of AI methods, several challenges remain for the real-world deployment, including issues related to data privacy, adaptability to dynamic conditions, and broader ethical considerations[129]. For instance, AI models may generate inaccurate or fabricated outputs (hallucinations) due to issues in training data quality, bias, and inconsistencies between training and deployment conditions [124]. Additionally, due to the data accessibility and AI’s black-box nature, they are vulnerable to attacks such as backdoor manipulation and data exploitation.
In addition, sensitive information such as customer consumption patterns and infrastructure data cannot be freely exchanged due to compliance requirements under regulatory frameworks such as the EU General Data Protection Regulation (GDPR) [36] and the North American Electric Reliability Corporation Critical Infrastructure Protection (NERC CIP)[76]. These constraints lead to fragmented data silos and reduced model effectiveness. To address these limitations, federated learning (FL) has emerged as a distributed machine learning framework that enables collaborative model training without sharing raw data. Instead, local models are trained at individual data sources and only model parameters are exchanged [131]. This approach improves scalability and supports learning across geographically distributed and heterogeneous power system environments. However, challenges such as communication overhead, non-identical data distributions, security vulnerabilities, and aggregation complexity can affect model performance [26].